OpenStudy (anonymous):

Solve -6x-2y=38-x-3y=9 A) (–6, 1) B) (–1, 6) C) (9, –6) D) (–6, –1)

OpenStudy (anonymous):

Solve for x or y and plug it back in to the respective variable.

OpenStudy (anonymous):

so first, you're going to put all the x on one side, and all the y on the other. Let's add 6x and 3y to both sides, which will gives you: y= 5x + 38 = 9 Now, subtract 9 from both sides, which gives you: y= 5x + 29 From that equation above, plug the points (x,y) in and see if the coordinates matches. 1 = 5(-6) + 29 WRONG! 6 = 5(-1) + 29 WRONG! -6 = 5(9) + 29 WRONG! -1 = 5(-6) + 29 CORRECT!